### 07 Mar 2024 v3

# IRConStyle: Image Restoration Framework Using Contrastive Learning and Style Transfer

Abstract
Recently, the contrastive learning paradigm has achieved remarkable success in high-level tasks such as classification, detection, and segmentation. However, in low-level tasks, like image restoration, contrastive learning is limited, and its effectiveness is uncertain. This raises a question: Why does the contrastive learning paradigm not yield remarkable results in image restoration? In this paper, we conduct in-depth analyses and propose three guidelines to address the above question. In addition, inspired by style transfer and based on contrastive learning, we propose a novel module for image restoration called ConStyle, which can be efficiently integrated into any U-Net structure network. By leveraging the flexibility of ConStyle, we develop a general restoration network for image restoration. The ConStyle and general restoration network together form an image restoration framework, namely IRConStyle. To demonstrate the capability and compatibility of ConStyle, we replace the general restoration network with transformer-based, CNN-based, and MLP-based networks, respectively. We perform extensive experiments on various image restoration tasks, including denoising, deblurring, deraining, and dehazing. The results on 19 benchmarks demonstrate that ConStyle can be integrated with any U-Net-based network and significantly enhance the performance of the network. For instance, ConStyle NAFNet significantly outperforms the original NAFNet on SOTS outdoor (dehazing) and Rain100H (deraining) datasets, with PSNR improvements of 4.16 dB and 3.58 dB with 85% fewer parameters. The code is available at: https://github.com/Dongqi-Fan/IRConStyle
1 Introduction
Figure 1: The style transfer experiment by StyleFlow [20] is performed to observe whether specific degradation of the degraded image could be transferred to a clean image.
Image Restoration (IR) aims to reconstruct degraded images into clean images, which can be applied in various fields such as surveillance, medical images, and autonomous driving. Recently, various IR works [39, 64, 37, 72, 57, 36, 49, 50, 26] have shown brilliant results and great prospects. However, these methods only focus on the specific module and operator design, ignoring the importance of overall network architecture and training strategy.
Typically, contrastive Learning (CL) paradigm [^1] In this paper, the CL paradigm refers to the use of contrastive loss, additional data structures to store positive or negative samples, and a large number of samples in calculating contrastive loss. [24, 12, 23, 16, 5, 10, 9] is considered a type of self-supervised or unsupervised learning, which has shown great potential in high-level tasks, often achieving or even surpassing the performance of supervised counterpart in downstream tasks. However, to the best of our knowledge, only two works [58, 32] have adopted the CL paradigm in the IR domain. Most studies [59, 19, 11, 46, 13, 14] only use the contrastive loss. Note that whether these works adopt the CL paradigm or not, the final effect is not expected. This raises a question: Why does the CL paradigm not achieve great success in IR? For this, we conduct in-depth thinking and propose three guidelines to fully leverage the potential of the CL paradigm in IR (see Sec. 3.1 for details): Firstly, additional data structures should be used to store enough positive/negative samples. Secondly, it is necessary to fully use the encoder’s intermediate feature map and the output latent code (the intermediate feature map and latent code referred to as latent feature). Finally, a reasonable pretext task should be adopted.
Image style transfer refers to transferring the style of a style image to a content image so that the content image retains its original content while also adopting the style of the style image. Since style transfer can transfer a specific image style, we wondered if we could treat degradations, such as blur and haze, as a kind of style and prevent the style image from transferring these degradations to the content image during IR. To test our idea, based on StyleFlow [20], we conduct a transfer experiment (Fig. 1) using the clean image as the content image and the degraded image as the style image. The noise, blur, and haze are treated as separate styles. Fig. 1 demonstrate that StyleFlow can effectively treat specific degradations in different degraded images as distinct styles and transfer these styles to the content image. Based on this observation, we utilize two simple loss functions, namely content loss and style loss, to guide the ConStyle in avoiding the transfer of such degradations to the latent feature. As depicted in Fig. 2 (c), the latent feature is encouraged to be far away from the degradation space and as close as possible to the clean space under the influence of content, style, and contrastive loss.
In this paper, we introduce a novel plug-and-play module for IR called ConStyle (Fig. 2 (a)), based on the guidelines mentioned above and style transfer. The ConStyle is built upon the MoCo [24] framework, replaces the original pretext task with a more suitable one for IR, and adopts a different encoder architecture. In addition, content loss and style loss are introduced to achieve reverse style transfer. Furthermore, thanks to the flexibility of ConStyle, we develop a general U-Net restoration network for IR. Together, ConStyle and the general restoration network (Fig. 2 (b)) form the IR framework: IRConStyle, with the module in the restoration network being replaceable by any specific operator. The choice of a U-Net structure for the restoration network is motivated by its ability to extract multi-scale information and its consistency with many existing IR networks [56, 60, 4, 42, 73, 69, 35, 17]. The latent feature extracted by ConStyle, which also contains multi-scale information, integrates well with this structure. Finally, to demonstrate the capability and compatibility of ConStyle, we replace the restoration network with Restormer [69] (a transformer-based network), NAFNet [8] (a CNN-based network), and MAXIM-1S [56] (an MLP-based network), called ConStyle Restormer, ConStyle NAFNet and ConStyle MAXIM-1S, and conduct extensive experiments on various IR tasks including denoising, deblurring, dehazing, and deraining. The extensive experiments show that ConStyle can integrate with different types of networks and perform reliably in different IR tasks.
Our contributions can be summarized as follows:
  • We thoroughly analyze the problems and limitations of CL in IR and propose three guidelines that can enhance the effectiveness of contrastive learning in IR.
  • We propose a general IR framework: IRConStyle consisting of ConStyle and a general restoration network. Notably, the module within the restoration network can be easily substituted with any specific operator.
  • We innovatively apply the idea of image style transfer to CL. Which significantly improves the performance of ConStyle without introducing any parameters or increasing the inference burden.
Figure 2: The overall architecture of IRConStyle ((a)(b)). (a) ConStyle. (b) General restoration network, where the Preprocess, Process, and Finetune modules can be replaced with any specific operator. (c) Under the influence of Con. (contrastive), content, and style loss, the latent feature keeps moving closer to clean space and away from degradation.
2 Related Work
2.1 Contrastive Learning
Contrastive learning, also known as self-supervised or unsupervised learning, is a method that enables efficient pre-training on large-scale unlabeled data. It achieves this by setting up ingenious pretext tasks. This method often performs as well as, or even surpasses, the supervised counterpart on certain tasks by fine-tuning on labeled data after pre-training [5, 10, 9, 48, 74, 3]. The InstDisc [62] proposes Instance Discrimination as a pretext task for the first time. In this approach, a batch of images is fed into an encoder at each iteration to obtain the corresponding latent code, which is then stored in a memory bank. The contrastive loss of each iteration is calculated between the current latent code and all historical latent codes in the memory bank; InvaSpread [67] also uses Instance Discrimination as a pretext task but does not utilize a memory bank or other data structure for storage. It solely relies on an encoder for end-to-end training, thus not fully realizing the potential of contrastive learning. To address this, MoCo [24] leverages the queue first-in first-out property and the strategy of exponential moving average (EMA) to construct data structures that enable effective contrastive learning. This approach has shown excellent performance in high-level tasks. However, most methods [34, 61, 63, 46, 13, 14] in IR simply adopt the contrastive loss of the CL paradigm. They do not utilize large and consistent data structures to store latent codes, making it challenging to fully leverage the potential of the CL paradigm, even with large batch sizes.
2.2 Image Restoration
In recent years, we have seen remarkable achievement in IR [71, 26, 54, 37, 28, 27, 41, 29, 6]. IPT [7] first adopts the transformer into low-level tasks to achieve image super-resolution, denoising, and deraining. At the same time, contrastive loss is adopted in IPT where patches within the same image are considered positive samples, while patches from different images are treated as negative samples. To reduce the computational complexity of the transformer, the sliding window mechanism [43, 39] is introduced based on ViT [18]; To further reduce the computation, [69] propose a transformer block specifically designed for IR; Additionally, [38, 15] utilize a mixed attention mechanism within the transformer block to extract more informative features; However, [8] argues that complex nonlinear operators and attention mechanisms may be unnecessary, as satisfactory results could be achieved using a simple CNN network. For [70], it proposes the Multi-Stage Progressive Restoration method, which employs a U-Net architecture for the first two stages. Inspired by the remarkable success of MLP [55, 40], [56] leverage gMLP [40] to extract both local and global information from the feature map. Different from them, we introduce a simple yet effective Image Restoration framework, in which the modules are replaceable by any specific operator.
2.3 Style Transfer
Image style transfer [20, 22, 68, 65] is a process that involves taking two input images, a content image and a style image, and extracting their respective content and style to generate a new image. This new image retains the content of the original content image while incorporating the style of the style image. CNN is adopted for the first time [22] to realize the style transfer, and then [68] implements style transfer through Wavelet Transforms. Since image style transfer is a generative task, Generative Adversarial Network is applied to realize the style transfer in an unsupervised manner [65]. In our training process, we leverage the style transfer to guide the encoder’s latent feature away from degradation space and towards clean space.
3 Method
In this section, we first provide a detailed analysis of why the CL paradigm is not effective in IR and draw corresponding conclusions. Subsequently, we introduce the entire IRConStyle. Finally, we demonstrate how ConStyle integrates with Restormer, NAFNet, and MAXIM-1S.
3.1 Analysis of Contrastive Learning in IR
Limitations of CL paradigm in IR. We believe that the main reason why CL is not as widely used in low-level tasks compared to high-level tasks is due to the difference in the output of the network. In high-level tasks, the network outputs an abstract semantic feature, while in low-level tasks, the output is an image of the same size as the input. Specifically, the network in high-level tasks does not have strict requirements on the size of the input image and can output an n-dimensional feature vector. However, in low-level tasks, the input and output of the network usually need to remain the same size or even require the output image to be several times larger than the input (e.g., image super-resolution). If a single network is used for the CL paradigm, it cannot guarantee that an image of the same size as the input can be restored from the corresponding latent code. This is because the network does not know the size of the real image in advance, and the real image size can even reach 4K, leading to the dimensions of the latent code being much lower than the dimensions of the real image. Therefore, most works in low-level tasks only use contrastive losses of the CL paradigm. Another solution is to employ two networks [58, 32], one for CL and one for restoration. The former is used to extract latent code and feed it to the latter, which utilizes such latent code to help rebuild the degraded image into a clean one. It is important to note that whether these works adopt the CL paradigm or not, the final effect is not significantly improved. Here, we propose three guidelines to fully leverage the CL paradigm in IR, and ablation studies are made in Tab. 7.
Figure 3: The comparison of CL part in different networks. (a) MoCo. (b) MoCo adopted in DASR and AirNet. (c) ConStyle. (d) The momentum encoder and encoder architecture in ConStyle. Where Con. stands for Contrastive.
Guideline 1. Additional data structures should be used to store enough positive/negative samples. For IR, the size of the feature map in the model usually remains unchanged from input to output. This poses a challenge when using large batch sizes, such as 4096 as seen in SimCLR [9] and BYOL [23]. The small batch size results in very few positive samples, negative samples, and queries participating in the contrastive loss calculation during each iteration. As a result, the network struggles to distinguish between positive samples, negative samples, and queries, leading to poor performance. For instance, RLFN [31], the winner of NTIRE 2022 [30], used a batch size and patch size of 64 in its training only with contrastive loss. In its ablation experiment, the PSNR only increased by 0.01 dB on Set5, Set14, and BSD100 compared to the baseline without contrastive loss. The same situation happens in [63], where the improvement over baseline is only 0.04 dB. Conclusion: For IR, to give full play to the role of contrastive learning and avoid the huge memory occupation brought by the large batch size, additional data structures should be used to store enough positive/negative samples. Following the MoCo, DASR and AirNet, we use the queue in our ConStyle (Fig. 3 (c)) to store the negative sample.
Guideline 2. It is necessary to make full use of the encoder’s latent feature. Even though DASR [58] and AirNet [32] adopt the complete CL paradigm (Fig. 3 (b)), the final result is not very satisfactory. This is because they do not fully utilize the latent feature of the encoder (i.e. the latent code and intermediate feature map). As mentioned earlier, both DASR and AirNet consist of a CL network and a restoration network. However, in the CL part, they only consider the highly characterized latent code as input to the restoration network. It ignores the non-highly characterized feature map, which contains hierarchical semantic information that can assist in both the restoration process and the understanding of the highly characterized latent code in the restoration network. Besides, one needs to be careful if the latent code is placed in an inappropriate position within the restoration network, as it becomes difficult to integrate it with the local feature. Conclusion: We should not only use the latent feature but also avoid the problem of feature semantic mismatch. Therefore, in our ConStyle, we not only utilize latent features but also incorporate them into different levels of the U-Net-style restoration network through the Affine Transform.
Guideline 3. A reasonable pretext task should be adopted. Another reason why DASR and AirNet perform poorly is that they use a suboptimal pretext task. In most CL paradigms, Instance Discrimination is used as a pretext task. Although this method is effective for downstream tasks like classification, segmentation, and detection, which require the network to recognize and extract semantic information. We believe it is not suitable for low-level tasks such as IR. In IR, what we need is the ability to remove degradation and reconstruct degraded images into clean ones, rather than the ability to distinguish and extract different semantic information from images. Therefore, Instance Discrimination should not be used as DASR and AirNet did (Fig. 3 (b)). Conclusion: Reasonable pretext tasks should be adopted in the IR domain. In ConStyle, we input clean images to the momentum encoder and set the output as positive samples, while degraded images are input to the encoder and set the output as queries. Then, the obtained queries are fed into the queue, which is then set as the negative samples.
Figure 4: The detailed structure of the original models (a)(b)(c) and the ConStyle models (d)(e)(f). DC represents the downsample and concat operation, and UC represents upsample and concat operation.
3.2 Image Restoration Framework: IRConStyle
IRConStyle (Fig. 2) consists of two main components: ConStyle and the general restoration network. ConStyle is responsible for extracting latent features and passing them to the restoration network, which is responsible for the restoration process. The training and inference flow of our IRConStyle can be described as follows:
$I_{restored}=G(E(I_{degraded},I_{clean}),I_{degraded})$ (1)
$I_{restored}=G(E(I_{degraded}),I_{degraded})$ (2)
Where $E$ is the ConStyle, $G$ is the general restoration network, $I_{degraded}$ is the degraded image, $I_{clean}$ is the clean image corresponding to $I_{degraded}$, and $I_{restored}$ is the restored clean image. In ConStyle, we employ both the encoder and momentum encoder in the training (as described in Eq. 1), and remove the momentum encoder in the inference (as described in Eq. 2). To ensure that the latent features are far from the degradation space and closer to the clean space, the content and style loss are introduced. The total loss in our IRConStyle is:
$L_{total}=L_{style}+L_{content}+L_{InfoNCE}+L_{1}$ (3)
$L_{1}$ loss is a commonly used loss function in the IR to measure the distance between $I_{clean}$ and $I_{restored}$. We will describe the style loss ($L_{style}$), content loss ($L_{content}$), and InfoNCE loss ($L_{InfoNCE}$) in detail in the next section.
3.3 ConStyle
Following MoCo, DASR, and AirNet, our encoder and momentum encoder (Fig. 3 (d)) is a siamese network consisting of several convolutions and an MLP. We also update the encoder by backpropagation and use EMA to slowly update the momentum encoder through the encoder. As analyzed before, the use of Instance Discrimination is detrimental to IR. In DASR and AirNet (Fig. 3 (b)), degraded images are fed into the momentum encoder and encoder separately, with the output $k$ representing the positive sample and the output $q$ representing the query. Each iteration adds the obtained $k$ to the queue, and then the entire queue is treated as negative samples. In contrast, in our ConStyle, we input clean images to the momentum encoder and set the output $k$ as the positive sample, while degraded images are fed into the encoder, and the output $q$ is set as the query. Each iteration adds the obtained $q$ to the queue, and then the entire queue is considered negative sample. Consistent with previous work [24, 12, 58, 32], we also use InfoNCE as our contrastive loss:
$L^{DASR,AirNet}_{InfoNCE}=-\log\frac{\exp({q{\cdot}k^{\prime}/t})}{\sum_{i=0}^{N}\exp({q{\cdot}Q_{i}/t})}$ (4)
$L^{ConStyle}_{InfoNCE}=-\log\frac{\exp({q{\cdot}k/t})}{\sum_{i=0}^{N}\exp({q{\cdot}Q_{i}/t})}$ (5)
Eq. 4 represents the $L_{InfoNCE}$ used in DASR and AirNet, while Eq. 5 represents the $L_{InfoNCE}$ used in ConStyle. In these equations, $t$ represents a fixed temperature hyper-parameter, $Q$ represents the queue, and $N$ represents the queue size. $k^{\prime}$ is obtained by inputting $I_{degraded}$ into the momentum encoder, $k$ is obtained by inputting $I_{clean}$ into the momentum encoder, and $q$ is obtained by inputting $I_{degraded}$ into the encoder. $\sum_{i=0}^{N}\exp({q{\cdot}Q_{i}/t})$ indicates the computation between query and negative samples in the queue at each iteration.
Method Down. 1 Down. 2 Down. 3 Down. 4 Up. 4 Up. 3 Up. 2 Up. 1 Total
Restormer [69] Params(K) 10.3 41.4 165.8 0 0 2654.2 663.5 165.8 3701.3
GFLOPs 0.339 0.339 0.339 0 0 1.358 1.358 1.358 5.091
ConStyle Restormer Params(K) 5.1 20.7 46.6 0 0 801.7 1368.5 764.9 3007.8
GFLOPs 0.169 0.169 0.095 0 0 0.495 3.056 6.715 10.699
NAFNet [8] Params(K) 32.8 131.3 524.8 2098.1 32.7 131.0 524.2 2097.1 5572.4
GFLOPs 0.268 0.268 0.268 0.268 0.267 0.267 0.268 0.268 2.142
ConStyle NAFNet Params(K) 5.1 20.7 46.6 0 0 801.7 1368.5 764.9 3007.8
GFLOPs 0.169 0.169 0.095 0 0 0.495 3.056 6.715 10.699
MAXIM-1S [56] Params(K) 16.4 65.6 262.2 0 0 65.6 32.8 8.2 450.9
GFLOPs 0.134 0.134 0.134 0 0 0.033 0.067 0.067 0.569
ConStyle MAXIM-1S Params(K) 2.3 9.2 20.7 0 0 331.7 331.7 147.4 843.2
GFLOPs 0.075 0.075 0.042 0 0 0.169 0.679 1.206 2.246
Table 1: Comparison of the parameters and computations generated by using different upsample (Up.) and downsample (Down.) operation for ConStyle models and original models.
Method Params (M) GFLOPs Speed (us)
Restormer [69] 26.12 70.49 60
ConStyle Restormer 15.57 74.92 61
NAFNet [8] 87.40 49.13 53
ConStyle NAFNet 12.74 46.97 54
MAXIM-1S [56] 8.17 21.58 53
ConStyle MAXIM-1S 8.10 25.58 52
Table 2: Model (w/wo ConStyle) parameters, computation, and inference speed comparison.
Following guideline 3, we not only input the latent code $q$ into the restoration network but also input the feature map generated in the encoder into it using the Affine Transform. Because the feature map in the encoder is not as highly characterized as the latent code q. If only highly characterized $q$ is fed into the restoration network, it is difficult for the latter to fuse $q$ with local features, causing the problem of feature semantic mismatch. Furthermore, we leverage the idea of style transfer to guide the encoder in producing latent features that are closer to the clean space and further away from the degradation space. Specifically, we set up content loss and style loss to ensure this process:
$L_{style}=-(\|q^{T}_{1}q_{1}-q^{T}q\|_{2}+\|q^{T}_{2}q_{2}-q^{T}q\|_{2})$ (6)
$L_{content}=\|k^{T}k-q^{T}q\|_{2}$ (7)
In each iteration, there are both incoming and outgoing latent codes in the queue. $q_{1}$ represents the latent code that is currently coming out of the queue, and $q_{2}$ represents the latent code that is about to come out of the queue. In [20, 22], VGG [53] and gram matrix are used to extract style and content features. Then MSE is used to calculate style loss and content loss. Instead, our ConStyle, with the assistance of highly characterized $k$ and $q$, directly computes the corresponding content loss and style loss through MSE. This approach brings significant benefits at a minor training cost and without any inference burden.
3.4 ConStyle with Other Network
To demonstrate the versatility of our IRConStyle, we replace the restoration network with other state-of-the-art networks: Restormer [69] (transformer-based), NAFNet [8] (CNN-based), and MAXIM-1S [56] (MLP-based) (Fig. 4). Restormer, NAFNet, and MAXIM-1S use different operators to implement the downsample and upsample modules. For instance, Restormer utilizes PixelUnShuffle for downsample and PixelShuffle for upsample, whereas MAXIM-1S uses stride Conv for downsample and ConvTranspose for upsample. In ConStyle Restormer, ConStyle NAFNet, and ConStyle MAXIM-1S, we use unified downsample and upsample modules of the proposed general restoration network, where the downsample consists of 1$\times$1 Conv and PixelUnShuffle, while the upsample consists of 1$\times$1 Conv and PixelShuffle.
To highlight the powerful capabilities of ConStyle and demonstrate that the improvement of the ConStyle models is not brought by simply expanding the scale of the network, we downscale ConStyle models by reducing the width or depth. Because of the introduction of ConStyle part, the parameters of networks will be increased by 1.19M. For instance, in NAFNet, the number of blocks in U-Net left, U-Net bottom, and U-Net right is [5, 6, 7, 7], 7, and [7, 7, 6, 5] respectively, with a network width of 64. In ConStyle NAFNet, the number of blocks in U-Net left, U-Net bottom, and U-Net right is [7, 8, 9], 9, and [9, 8, 7] respectively, with a network width of 48. Despite the additional 1.19M parameters introduced by ConStyle, our ConStyle NAFNet has 74.66M (85%) fewer than NAFNet.
4 Experiments
Method CBSD68(***) [45] Urban 100(***) [25] CBSD68 [45] Urban 100 [25]
$\sigma$=15 $\sigma$=25 $\sigma$=50 $\sigma$=15 $\sigma$=25 $\sigma$=50 $\sigma$=15 $\sigma$=25 $\sigma$=50 $\sigma$=15 $\sigma$=25 $\sigma$=50
Restormer [69] 34.32 31.71 28.51 34.87 32.65 29.63 34.37 31.74 28.53 35.00 32.75 29.71
ConStyle Restormer 34.33 31.71 28.51 34.89 32.66 29.64 34.37 31.74 28.52 35.01 32.74 29.71
Differ. +0.01 0 0 +0.02 +0.01 +0.01 0 +0 +0 -0.01 -0.01 +0
NAFNet [8] 34.26 31.64 28.46 34.62 32.36 29.29 34.32 31.71 28.52 34.80 32.65 29.63
ConStyle NAFNet 34.31 31.69 28.50 34.82 32.58 29.55 34.34 31.71 28.52 34.91 32.64 29.63
Differ. +0.05 +0.05 +0.04 +0.20 +0.22 +0.26 +0.02 0 +0 +0.11 -0.01 +0
MAXIM-1S [56] 34.25 31.63 28.43 34.52 32.23 29.10 34.28 31.65 28.44 34.63 32.31 29.14
ConStyle MAXIM-1S 34.25 31.63 28.43 34.56 32.26 29.10 34.28 31.65 28.44 34.64 32.32 29.13
Differ. 0 0 0 +0.04 +0.03 0 0 0 0 +0.01 +0.01 -0.01
Table 3: RGB image denoising tested by PSNR. Where (***) indicates models are trained with random sigma (0 to 50), while other indicates models are trained with fixed sigma. Blue means better and red means worse.
Method GoPro [47] RealBlur-R [51] RealBlur-J [51] HIDE [52]
Restormer [69] 31.28/0.9188 33.92/0.9446 26.61/0.8286 30.16/0.9097
ConStyle Restormer 31.45/0.9208 33.94/0.9454 26.63/0.8288 30.02/0.9098
Differ. +0.17/ +0.0020 +0.02/ +0.0008 +0.02/ +0.0002 +0.04/ +0.0001
NAFNet [8] 31.29/0.9193 33.75/0.9434 26.62/0.8219 30.11/0.9095
ConStyle NAFNet 31.56/0.9230 33.89/0.9441 26.61/0.8308 30.24/0.9106
Differ. +0.27/ +0.0037 +0.14/ +0.0007 -0.01/ +0.0089 +0.13/ +0.0011
MAXIM-1S [56] 30.50/0.9039 33.89/0.9431 26.48/0.8230 29.22/0.8918
ConStyle MAXIM-1S 30.77/0.9093 33.93/0.9435 26.54/0.8261 29.26/0.8951
Differ. +0.27/ +0.0054 +0.04/ +0.0004 +0.06/ +0.0031 +0.04/ +0.0033
Table 4: Image deblurring tested by PSNR/SSIM. Blue means better and red means worse.
4.1 Experimental Settings
[…]
Following the Restormer, NAFNet, and MAXIM-1S, for deblurring, we train all models using GoPro [47] and evaluate them on the GoPro, RealBlur-R [51], RealBlur-J [51], and HIDE [52]; For denoising, DFWB (DIV2K [1], Flickr2K, WED [44], and BSD500 [2]) are used for training, and CBSD68 [45] and Urban100 [25] are used for evaluation; For dehazing, OTS [33] is used for training and SOTS [33] outdoor is for evaluation; For deraining, Rain14000 [21] and Rain100H [66] are used for training and Rain100H and Rain100L [66] are for evaluation.
4.2 Model Analyses
Tab. 2 presents a comparison of parameters, computation, and speed between the original models and our ConStyle models. All the results are obtained using input data of size (2,3,128,128), and the speed is the average of 10,000 inference. As shown in Tab. 2, despite the introduction of additional parameters (1.19M), ConStyle models have fewer parameters than the original models. While the GFLOPs may not accurately measure the inference burden, we provide the inference time. The decrease in the number of parameters and increase in computation may seem counterintuitive. For instance, ConStyle Restormer reduces the number of parameters by 10.55 M but increases the computation by about 3 GFLOPs. This discrepancy arises from the use of different operators for downsampling and upsampling. As discussed in Sec. 3.4, in our ConStyle models, we replace the original models’ downsampling and upsampling with those from the restoration network. Tab. 1 demonstrates that this replacement adds 5.64, 8.56, and 1.67 GFLOPs to ConStyle Restormer, ConStyle NAFNet, and ConStyle MAXIM-1S, respectively.
Figure 5: The comparison of deblurring, dehazing, denoising, and deraining between ConStyle models and original models.
4.3 Experiment Results
Image Deblurring. ConStyle NAFNet, ConStyle MAXIM-1S, NAFNet, and MAXIM-1S are trained for 100 epochs, while ConStyle Restormer and Restormer are for 75 epochs. As shown in Tab. 4, except for the slightly worse performance of ConStyle NAFNet on the RealBlur-J and ConStyle MAXIM-1S on the RealBlur-R, in general, our ConStyle models can improve the performance of original models with the help of ConStyle. For example, ConStyle NAFNet significantly outperforms NAFNet on GoPro, RealBlur-R, and HIDE, with improvements of 0.27 dB, 0.14 dB, and 0.13 dB on PSNR, respectively. ConStyle Restormer also significantly outperforms Restormer on the GoPro, with a PSNR increase of 0.17 dB.
Image Denosing. ConStyle NAFNet, ConStyle MAXIM-1S, NAFNet, and MAXIM-1S are trained for 100 epochs, while ConStyle Restormer and Restormer are for 75 epochs. Tab. 3 demonstrates that in the random sigma benchmarks, the PSNR of all ConStyle models is higher than or equal to the corresponding original models. It’s worth noting that, with 85% fewer parameters, ConStyle NAFNet significantly outperforms NAFNet on the Urban100 (+0.20, +0.22, and +0.26 dB).
Image Dehazing. ConStyle NAFNet, ConStyle MAXIM-1S, NAFNet, and MAXIM-1S are trained for 21 epochs, while ConStyle Restormer and Restormer are trained for 16 epochs. Tab. 5 demonstrates that ConStyle Restormer performs slightly better than Restormer, while ConStyle MAXIM-1S performs slightly worse than MAXIM-1S. Notably, ConStyle NAFNet achieves a significant performance boost of 4.16 dB compared to NAFNet.
Image Deraining. ConStyle NAFNet, ConStyle MAXIM-1S, NAFNet, and MAXIM-1S are trained for 46 epochs, while ConStyle Restormer and Restormer are trained for 35 epochs. Tab. 5 demonstrates that except that ConStyle MAXIM-1S performs worse than MAXIM-1S on the Rain100H, all ConStyle models significantly outperform the original models on all benchmarks.
Method SOTS outdoor [33] Rain100H [66] Rain100L [66]
Restormer [69] 30.83/0.9761 30.46/0.9047 37.82/0.9832
ConStyle Restormer 30.85/0.9760 30.82/0.9095 38.23/0.9844
Differ. +0.02/ -0.0001 +0.36/ +0.0038 +0.41/ +0.0012
NAFNet [8] 26.57/0.9513 26.94/0.8455 34.97/0.9685
ConStyle NAFNet 30.73/0.9743 30.52/0.9061 37.79/0.9832
Differ. +4.16/ +0.0230 +3.58/ +0.0606 +2.82/ +0.0147
MAXIM-1S [56] 30.47/0.9736 29.50/0.8872 35.54/0.9717
ConStyle MAXIM-1S 30.59/0.9713 29.20/0.8845 36.23/0.9748
Differ. +0.12/ -0.0023 -0.30/ -0.0027 +0.69/ +0.0031
Table 5: Image dehazing and deraining tested by PSNR/SSIM. Blue means better and red means worse. The first column represents dehazing, and the others represent deraining.
Content Style InfoNCE PSNR Content Style InfoNCE PSNR
31.45 32.36
29.98 31.61
31.50 29.70
31.56 32.58
26.94 26.57
27.60 28.29
27.51 29.87
30.52 30.73
Table 6: Ablation studies on ConStyle NAFNet to evaluate the performance of loss functions. The table in the top left, top right, bottom left, and bottom right represent the results of deblurring on GoPro [47], denoising on Urban100 [25] ($\sigma$=25), deraining on Rain100H [66], and dehazing on SOTS outdoors [33], respectively.
Baseline Guideline 1 Guideline 2 Guideline 3
PSNR 31.56 28.82 31.33 31.45
Table 7: Ablation studies of ConStyle NAFNet on GoPro [47] to evaluate proposed three guidelines.
4.4 Ablation Studies
We conduct two ablation experiments to evaluate the performance of loss functions (Tab. 6) and proposed three guidelines (Tab. 7). In Tab. 7, our ConStyle NAFNet is set as a baseline, in which the content, style, and InfoNCE loss are used as well as the latent feature map. For guideline 1, we change the Queue size of the ConStyle from 65760 to 16; For guideline 2, the latent feature maps are removed. For guideline3, we move the Queue from behind the Encoder back to behind the Momentum Encoder, in line with AirNet [32], DASR [58], and MoCo [24].
5 Conclusion
In this paper, we analyze the limitations of the CL paradigm in IR and propose three guidelines to fully leverage CL in IR. In addition, we introduce an image restoration framework: IRConStyle, consisting of ConStyle and the restoration network. ConStyle extracts latent features and passes them to the restoration network which is responsible for image restoration.